# Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents

### v1

@@ 4.2 Green citations
Our baseline exercise is reported in column (4) of Table 4. We find that true greens are cited about $1\%$ less than the rest of the green patents, as from classifications in previous economic literature. Please note how the results change when we switch from column (2) to column (3) after we include fixed effects for technological classes and priority years. One possible explanation for the negative premium on citations by true green patents is that environmental technologies are relatively newer than brown technologies; therefore, there is less scope for citations by following novel inventions, which introduce other environmental technologies by relying on scarcer knowledge about environmental protection.
@@
Adj. R squared 0.001 0.160 0.378 0.384
N. obs. 4,049,540 4,049,540 4,049,540 4,049,540
  • Note: The table reports results after dummy regressions where the dependent variable is the logarithm of the patent citations plus one. We consider the entire sample of green patents as considered by previous literature, and our coefficient of interest is the one on True green patents, after following our NLP approach. Patent controls include age and the number of patent family members. Errors are clustered at the patent family level. *** p$<$0.01, ** p$<$0.05, * p$<$0.1.
@@ 5.1 Firm-level premia
N. firms with green grants 3,137 3,137 2,407 3,111 3,137 2,576 1,133
N.obs. 11,320,979 11,320,979 9,337,060 11,253,885 11,320,984 8,997,862 1,784,972
Adj. R-squared 0.938 0.387 0.951 0.222 0.420 0.187 0.870
  • Note: The table reports results after dummy regressions where dependent variables are a chosen dashboard of firms’ outcomes, and the coefficients of interest are True green and Patenting. Firm size is measured by the number of employees, and industries are 2-digit NACE rev. 2. Sales, labor productivity, capital intensity, and TFP are expressed in logarithms. ROCE can have zero values and is transformed with a hyperbolic sine as proposed by Bellemare and Wichman (2020). Market share and EBIT are expressed in levels. Firm-level clustered standard errors. *** p$<$0.01, ** p$<$0.05, * p$<$0.1.
In particular, we observe that, after controlling for our NLP approach, firms that obtain true green show positive premia in market shares, labor productivity, and capital intensity. Interestingly, we record negative premia on both measures of profitability, the ROCE and the EBIT, while the coefficient of TFP for manufacturing firms is not statistically significant. Please note how important it is to control for Patenting. Regardless of the technology, firms that obtain a patent grant show almost always higher outcomes, with the notable exception of profitability measures.
At this stage, we cannot interpret the coefficients in Table 5 as the effect of patenting activity on firms’ competitiveness. We can only conclude that, apparently, firms with true green grants are statistically different from the ones with grants in other technological classes. In the next Section, we will provide a few additional elements to comment on thier statistical differences.
@@ 5.2 Green patenting, firms’ outcomes, and reverse causality
Table 6 shows the distribution of treated firms in our panel. The number of newly treated individuals ranges from 372 to 241 per year. Please note that the group of firms that have never obtained a true green grant is huge, and that is the set from which we need to obtain a counterfactual after our PSM.
@@
Year True green
2010 329
2011 372
2012 359
2013 347
2014 352
2015 327
2016 335
2017 369
2018 348
2019 373
2020 353
2021 318
2022 241
  • Note: The table reports the panel distribution of companies in our sample that had at least one true green grant, i.e., following our NLP classification, in our period of analysis. Please note that repetitions may occur if the same firm had a green patent granted in different years, this definition of treatment is in line with our cross-section application of the PSM model.
@@
After the match of each firm with at least one true green grant with its nearest-neighbour that did not, we can estimate the Average Treatment Effect on the Treated (ATET) by comparing the means of the two distributions. In Table 7, we find that statistically significant ATETs emerge primarily for sales, market shares, and labor productivity. If a company obtains a true green grant, it can rely on a remarkable increase in sales ($e^{1.357-1}\approx$ 143%), market share ($e^{0.033-1}\approx$ 38%), and labor productivity ($e^{0.085-1}\approx$ 40%). These effects align with the correlations observed in previous paragraphs. The main departure lies in the coefficients of capital intensity, ROCE, EBIT, and TFP. Profitability measures are not statistically significant after we introduce a matching strategy, while they correlate negatively with true green grants. The most likely explanation is that, after investing in green technologies, firms are, on average, at a loss. They still need to break even after the sunk cost of the R&D needed to obtain those grants. Yet, after we construct our counterfactual, there is no systematic difference in profitability between a company that invests in green technologies and a similar one that invests in any other technology.
@@
ATET 1.357*** 0.033*** 0.085*** -0.01 0.048 0.11 0.002
std. error 0.089 0.007 0.028 0.027 0.073 0.418 0.156
t-stat. 15.232 4.782 3.053 -0.372 0.655 0.263 0.014
Treated obs. 2,615 2,276 2,034 2,276 2,276 2,076 968
Untreated obs. 46,345 41,276 36,378 41,276 41,276 37,432 24,992
  • Note: The table reports ATTs computed via PSM. The treated companies have at least one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1.
@@ 6 Robustness and sensitivity checks
In this section we present robustness and sensitivity controls on previous results using our measures. The inclusion of the measure following Arts et al. (2021) reveals that the positive effects of green patenting are highly contingent on the novelty they introduce in their technology classes. Our NLP approach, used together with the NLP approach for novelty (8), suggests that green patenting delivers stronger economic returns when it is genuinely transformative.
@@
ATET 0.885*** 0.026 -0.014 -0.086 -0.047 0.409 0.403
std. error 0.243 0.018 0.067 0.058 0.142 0.875 0.464
t-stat. 3.643 1.433 -0.217 -1.481 -0.33 0.468 0.87
Treated obs. 530 499 440 499 499 467 157
Untreated obs. 11,135 10,421 9,331 10,421 10,421 9,841 5,724
  • Note: The table reports ATETs computed after a propensity score matching. The treated unit is a company that has been granted at least one high-novelty true-green patent for a single year, and it is matched with companies that have at least one high-novelty grant, regardless of the technology nature, in the period 2010-2022. High novelty is defined following the NLP approach by Arts et al. (2021), while true greens are defined following our NLP approach. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
As a first robustness check, we estimate the ATETs separately for small and large firms (Tables 10 and 9). Firms are classified based on the median number of employees [^9] Measured as the average number of employees over the 2010–2022 sample period in our sample, which is 3.6. For larger firms, the results are largely consistent with the baseline estimates. In contrast, for smaller firms, no statistically significant effects are observed across most outcomes, with the exception of a 14% positive effect on EBIT. This result may suggest that while smaller firms do not immediately experience growth in scale (e.g., sales or market share), they may achieve gains in operating efficiency or cost reductions from engaging in green patenting activities. In contrast, larger firms may already benefit from economies of scale and more established innovation pipelines, making the marginal benefit of green patenting less pronounced in profitability terms.
@@
ATET 1.382*** 0.032*** 0.093*** -0.046* 0.07 0.058 0.006
std. error 0.085 0.007 0.027 0.027 0.072 0.394 0.158
t-stat. 16.198 4.634 3.465 -1.727 0.967 0.148 0.036
p-value 0.0 0.0 0.001 0.084 0.333 0.883 0.971
Treated obs. 2,573 2,247 2,009 2,247 2,247 2,050 962
Untreated obs. 45,124 40,313 35,729 40,313 40,313 36,622 24,784
  • Note: The table reports ATETs computed after a propensity score matching only for firms with an average size (determined by number of employees) during the sample period, above the median value. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
@@
ATET -0.353 0.0 -0.078 -0.235 0.313 14.714** 0.187
std. error 0.72 0.003 0.509 0.452 0.881 7.349 0.217
t-stat. -0.49 0.057 -0.154 -0.519 0.355 2.002 0.862
p-value 0.624 0.954 0.878 0.604 0.723 0.046 0.395
Treated obs. 42 28 25 28 28 25 6
Untreated obs. 727 508 366 508 508 440 29
  • Note: The table reports ATETs computed after a propensity score matching only for firms with an average size (determined by number of employees) during the sample period, below or equal the median value. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
A second robustness check involves distinguishing between firms located in “old” and “new” EU member states (Tables 11 and 12). Specifically, we define old members as countries that joined the European Union before 2000, and new members as those that joined thereafter. No significant ATETs are found for firms in the new member states. However, for firms in the old member states, we observe a positive and statistically significant effect on both sales ($e^{0.803-1}\approx$ 82%) and market share ($e^{0.05-1}\approx$ 39%). One possible interpretation is that firms in old EU member states are better positioned to capitalize on green patenting activities due to more mature institutional environments, stronger regulatory frameworks, and better access to green innovation subsidies or infrastructure. In contrast, firms in new member states may still be in the process of integrating environmental standards and innovation systems, limiting their ability to convert green patenting into immediate competitive or market advantages. Alternatively, differences in domestic demand for green products and varying levels of international exposure could also contribute to the observed heterogeneity in outcomes.
Table 11: Average Treated Effects on the Treated after a propensity score matching - true-green patent and firms outcomes for first countries joining EU.
@@
ATET 0.803*** 0.05** 0.035 -0.02 0.027 0.039 -0.279
std. error 0.261 0.02 0.054 0.056 0.165 0.88 0.39
t-stat. 3.078 2.544 0.652 -0.367 0.163 0.044 -0.715
Treated obs. 457 432 378 432 432 403 137
Untreated obs. 1,724 1,579 1,421 1,579 1,579 1,424 720
  • Note: The table reports ATETs computed after a propensity score matching only for firms of countries joining the EU before the year 2000, i.e. Austria, Belgium, Denmark, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Portugal, Spain and Sweden. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
Table 12: Average Treated Effects on the Treated after a propensity score matching - true-green patent and firms outcomes for last countries joining EU.
Sales Mkt share Lab. prod. K-intensity ROCE EBIT TFP
ATET -0.141 0.043 0.119 -0.071 -0.177 3.31 -2.377
std. error 0.454 0.044 0.132 0.157 0.366 2.889 3.497
t-stat. -0.31 0.965 0.908 -0.452 -0.483 1.146 -0.68
Treated obs. 61 54 51 54 54 51 14
Untreated obs. 247 115 101 115 115 110 37
  • Note: The table reports ATETs computed after a propensity score matching only for firms of countries joining the EU after the year 2000, i.e. Bulgaria, Croatia, Cyprus, Czechia, Estonia, Hungary, Latvia, Lithuania, Malta, Poland, Romania, Slovakia and Slovenia. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
@@ 7 Conclusions and final remarks
This study’s primary contribution is the development and application of an NLP methodology for classifying “true” green patents. Our research suggests that this custom NLP approach can refine the identification of environmentally relevant patents compared to traditional, broader classification systems. Standard methods may risk conflating genuinely environmental technologies with conventional technologies. Our validation procedure, comparing the NLP approach against more conventional classifications, indicates that standard methods broadly overestimate green patents. We find that about 20% of those otherwise flagged as green according to standards in literature are actually true-green patents when we apply our text-analysis methodology. We show, however, an important degree of heterogeneity once we scroll down the different technological classes that are touched upon by true environmental technologies.
@@ A.1. Green-patents vocabulary
This section contains reference to our custom dictionary constructed for true green patents classification. Table A1 contains baseline expressions related to green patents and technologies produced by Chat GPT and Gemini and used to feed the Word2Vec model for the production of an enlarged dictionary. Table A2, A3, and A4 represent our final enlarged vocabulary for regex match into patents descriptive texts. It is made of two kinds of expressions:

### v2

@@ 4.3 Green citations
Our baseline exercise is reported in column (4) of Table 3. We find that true greens are cited about $4\%$ less than the rest of the green patents, as from classifications in previous economic literature. Please note how the results change when we switch from column (2) to column (3) after we include fixed effects for technological classes and priority years. One possible explanation for the negative premium on citations by true green patents is that environmental technologies are relatively newer than brown technologies; therefore, there is less scope for citations by following novel inventions, which introduce other environmental technologies by relying on scarcer knowledge about environmental protection.
@@
Adj. R squared 0.001 0.230 0.298 0.307
N. obs. 4,049,512 4,049,512 4,049,512 4,049,512
  • Note: The table reports results after dummy regressions where the dependent variable is the logarithm of the patent citations plus one. We consider the entire sample of green patents as considered by previous literature, and our coefficient of interest is the one on True green patents, after following our NLP approach. Patent controls include age and the number of patent family members. Errors are clustered at the patent family level. *** p<<0.01, ** p<<0.05, * p<<0.1.
@@ 5.1 Firm-level premia
N. firms with green grants 692 692 692 692 692 692 169
N. obs. 7,455,164 7,455,164 7,455,164 7,455,164 7,455,164 7,455,164 1,317,731
Adj. R-squared 0.947 0.454 0.952 0.224 0.469 0.188 0.873
  • Note: The table reports results after dummy regressions where dependent variables are a chosen dashboard of firms’ outcomes, and the coefficients of interest are True green and Patenting. Firm size is measured by the number of employees, and industries are 2-digit NACE rev. 2. Sales, labor productivity, capital intensity, and TFP are expressed in logarithms. ROCE can have zero values and is transformed with a hyperbolic sine as proposed by Bellemare and Wichman (2020). Market share and EBIT are expressed in levels. Firm-level clustered standard errors. *** p<<0.01, ** p<<0.05, * p<<0.1.
In particular, we observe that, after controlling for our NLP approach, firms that obtain true green show positive premia in sales, market shares, labor productivity, and capital intensity. Interestingly, we record negative premia on both measures of profitability, the ROCE and the EBIT, while the coefficient of TFP for manufacturing firms is not statistically significant. Please note how important it is to control for Patenting. Regardless of the technology, firms that obtain a patent grant show almost always higher outcomes, with the notable exception of profitability measures.
At this stage, we cannot interpret the coefficients in Table 4 as the effect of patenting activity on firms’ competitiveness. We can only conclude that, apparently, firms with true green grants are statistically different from the ones with grants in other technological classes. In the next section, we will provide a few additional elements to comment on their statistical differences.
@@ 5.2 Green patenting, firms’ outcomes, and reverse causality
Table 5 shows the distribution of treated firms in our panel. The number of newly treated individuals ranges from 326 to 225 per year.
@@
Year Treated firms Untreated firms
2010 283 7,715
2011 326 7,859
2012 322 6,907
2013 325 6,045
2014 322 6,426
2015 287 6,402
2016 278 6,751
2017 291 7,109
2018 291 6,773
2019 330 6,519
2020 296 6,402
2021 286 5,693
2022 225 4,675
  • Note: The table reports the panel distribution of companies in our sample that had at least one true green grant (treated), i.e., following our NLP classification, and no true green but at least one granted patent (untreated), in our period of analysis. Please note that repetitions may occur if the same firm had a green patent granted in different years, this definition of treatment is in line with our cross-section application of the PSM model.
@@
After the match of each firm with at least one true green grant with its nearest-neighbour that did not issue any, we can estimate the Average Treatment Effect on the Treated (ATET) by comparing the means of the two distributions. In Table 6, we find that statistically significant ATETs emerge primarily for sales, market shares, and labor productivity. If a company obtains a true green grant, it can rely on a remarkable increase in sales ($e^{1.318-1}\approx$ 137%), market share ($e^{0.036-1}\approx$ 38%), and labor productivity ($e^{0.113-1}\approx$ 41%). These effects align with the correlations observed in previous paragraphs. The main departure lies in the coefficients of capital intensity, ROCE, EBIT, and TFP. Profitability measures are not statistically significant after we introduce a matching strategy, while they correlate negatively with true green grants. The most likely explanation is that, after investing in green technologies, firms are, on average, at a loss. They still need to break even after the sunk cost of the R&D needed to obtain those grants. Yet, after we construct our counterfactual, there is no systematic difference in profitability between a company that invests in green technologies and a similar one that invests in any other technology.
@@
ATET 1.318*** 0.036*** 0.113*** -0.023 0.061 -0.125 -0.077
std. error 0.105 0.008 0.032 0.026 0.082 0.443 0.166
t-stat. 12.581 4.452 3.593 -0.859 0.748 -0.281 -0.463
Treated obs. 1,608 1,608 1,608 1,608 1,608 1,608 753
Untreated obs. 33,071 33,071 33,071 33,071 33,071 33,071 21,576
  • Note: The table reports ATTs computed via PSM. The treated companies have at least one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p<<0.01, ** p<<0.05, * p<<0.1.
@@ 6 Robustness and sensitivity checks
In this section we present robustness and sensitivity controls on previous results on the impact of green patenting activity on firm-level outcomes. The inclusion of the measure following Arts et al. (2021) reveals that the positive effects of green patenting are highly contingent on the novelty they introduce in their technology classes. Our NLP approach, used together with the NLP approach for novelty (7), suggests that green patenting do not generally delivers stronger economic returns when it is genuinely transformative. In fact, even if sales and market shares follow a high novelty true green grant with an increase of 108% ($e^{1.073-1}$) and 38% ($e^{0.036-1}$) respectively, the effect on profitability proxied by EBIT is larger negative. However, this result can be interpreted in light of a reasonable timeline for the innovation process, where intensive investments in R&D — particularly for highly innovative patented technologies — take time to generate returns and/or to absorb the impact of development costs on the balance sheet.
@@
ATET 1.073*** 0.036* 0.006 -0.025 -0.128 -1.581* 0.256
std. error 0.27 0.021 0.071 0.051 0.16 0.91 0.465
t-stat. 3.976 1.71 0.085 -0.483 -0.799 -1.736 0.551
Treated obs. 374 374 374 374 374 374 132
Untreated obs. 8,812 8,812 8,812 8,812 8,812 8,812 5,442
  • Note: The table reports ATETs computed after a propensity score matching. The treated unit is a company that has been granted at least one high-novelty true-green patent for a single year, and it is matched with companies that have at least one high-novelty grant, regardless of the technology nature, in the period 2010-2022. High novelty is defined following the NLP approach by Arts et al. (2021), while true greens are defined following our NLP approach. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p<<0.01, ** p<<0.05, * p<<0.1
As a first robustness check, we estimate the ATETs separately for small and large firms (Tables 8 and 9). Firms are classified based on the median number of employees [^9] Measured as the average number of employees over the 2010–2022 sample period in our sample, which is 214. For larger firms, the results are largely consistent with the baseline estimates, except for a negative effect on capital intensity. In contrast, for smaller firms, no statistically significant effects are observed across most outcomes.
@@
ATET 1.260*** 0.031*** 0.093*** -0.045* 0.071 -1.136*** 0.015
std. error 0.086 0.01 0.029 0.025 0.091 0.421 0.176
t-stat. 14.599 2.965 3.265 -1.781 0.786 -2.697 0.087
Treated obs. 1,191 1,191 1,191 1,191 1,191 1,191 582
Untreated obs. 15,957 15,957 15,957 15,957 15,957 15,957 10,619
  • Note: The table reports ATETs computed after a propensity score matching only for firms with an average size (determined by number of employees) during the sample period, above the median value. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p<<0.01, ** p<<0.05, * p<<0.1
@@
ATET -0.044 0.006 0.011 0.046 -0.034 0.5 -0.305
std. error 0.166 0.008 0.073 0.069 0.178 1.148 0.411
t-stat. -0.265 0.765 0.146 0.67 -0.191 0.436 -0.742
Treated obs. 416 416 416 416 416 416 171
Untreated obs. 15,634 15,634 15,634 15,634 15,634 15,634 9,861
  • Note: The table reports ATETs computed after a propensity score matching only for firms with an average size (determined by number of employees) during the sample period, below or equal the median value. The treated unit is a company that one true-green patent following our NLP approach in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p<<0.01, ** p<<0.05, * p<<0.1
Finally we applied our PSM analysis on the subset of green patents identified with our NLP methodology among all granted applications [^10] See Section 4.2. (not just the green as defined by standard algorithms). The results reported in Table 10 align with our baseline findings. This consistency further suggest a certain degree of coherency of our classification framework in accurately identifying green innovations, further signaling a certain rate of Type II errors inherent in standard classification systems.
Table 10: Average Treated Effects on the Treated after a propensity score matching - true-green patent classifications applied to all granted applications.
@@
ATET 1.173*** 0.022*** 0.065*** -0.006 0.068 -0.194 -0.088
std. error 0.073 0.005 0.022 0.019 0.059 0.311 0.12
t-stat. 15.966 4.24 2.914 -0.337 1.151 -0.624 -0.738
Treated obs. 2,927 2,927 2,927 2,927 2,927 2,927 1,573
Untreated obs. 32,537 32,537 32,537 32,537 32,537 32,537 21,613
  • Note: The table reports ATETs computed after a propensity score matching. The treated unit is a company that one true-green patent following our NLP approach applied not only to green patents (standard definition), but to all granted applications, in a single year, and it is matched with companies that have been granted at least one patent in 2010-2022, regardless of the technology nature. Propensity score matching is performed by nearest-neighbour without replacement, assuming a common support on the covariates. *** p<<0.01, ** p<<0.05, * p<<0.1
@@ 7 Conclusions and final remarks
This study’s primary contribution is the development and application of an NLP methodology for classifying “true” green patents. Our research suggests that this custom NLP approach can refine the identification of environmentally relevant patents compared to traditional, broader classification systems. Standard methods may risk conflating genuinely environmental technologies with conventional technologies. Our validation procedure, comparing the NLP approach against more conventional classifications, indicates that standard methods broadly overestimate green patents. We find that about 18.5% of those otherwise flagged as green according to standards in literature are actually true-green patents when we apply our text-analysis methodology. We show, however, an important degree of heterogeneity once we scroll down the different technological classes that are touched upon by true environmental technologies.
In addition, our analysis reveals the possibility of systematic Type I and Type II errors in traditional green patent classification that distort understanding of environmental innovation. The finding that only 18.5% of traditionally classified patents qualify as “true green” indicates substantial false positive rates, while identifying over 1 million overlooked patents ($\simeq 2.9\%$ during our sample period) demonstrates extensive false negatives. This kind of miss-classification is rooted in a limited or biased selection procedure carried on by expert during the definition of classical green technologies classification frameworks. By minimizing both error types, our methodology provides more reliable indicators for policy evaluation and resource allocation toward genuinely impactful green technologies.
@@ A.1. Green-patents vocabulary
This section contains reference to our custom dictionary constructed for true green patents classification. Table A1 contains baseline expressions related to green patents and technologies produced by Chat GPT and Gemini and used to feed the Word2Vec model for the production of an enlarged dictionary.
Table A1: Baseline key-expressions.
Baseline keywords
air purification eco friendly green design recycled packaging
alternative energy eco innovation green innovation recycled products
anaerobic digestion eco solutions green manufacturing regenerative braking
bio-based materials ecodesign green technology renewable energy
biodegradable polymers ecological design green transportation resource efficient
biodegradable products ecological footprint greenhouse gas resource recovery
biodiversity electric motors greenhouse gases solar geoengineering
biofuel electric vehicles hybrid vehicles solar panels
biomass energy conservation hydroelectric dams sustainability metrics
bioplastics energy efficiency hydroponics sustainability practices
carbon capture energy efficient hydropower sustainable agriculture
carbon filters energy optimization leed certified sustainable development
carbon neutral energy recovery life cycle sustainable farming
carbon offsetting energy savings lightweight materials sustainable forestry
carbon sequestration environmental certification low emission sustainable manufacturing
cellulosic plastics environmental compliance low impact sustainable materials
circular economy environmental conservation material compatibility sustainable practices
clean energy environmental labeling organic farming vertical farming
clean production environmental management photocatalysis waste minimization
clean technology environmental policy plant based waste recycling
climate adaptation environmental stewardship pollution control waste reduction
climate friendly fuel cells precision agriculture water conservation
compostable materials geothermal product lifespan water efficient
composting geothermal power public transportation wind turbines
dioxide capture global warming rainwater harvesting zero emissions
eco conscious green buildings recyclable materials zero waste
eco efficient green chemistry recycled materials
Baseline expressions in addition to the ones listed in Table, A2, and A3 represent our final enlarged vocabulary for regex match into patents descriptive texts. It is made of two kinds of expressions: